OpenStudy (anonymous):

the slant height of a cone is 15 cm and the radius of the base is 9 find the lateral area and the volume helpppp!!!!

OpenStudy (anonymous):

if i answer u give me a medal?

OpenStudy (anonymous):

yea if itz good

OpenStudy (anonymous):

T Y

OpenStudy (anonymous):

A= pi*R*L A= 9*pi*15 A= 424cm V=(pi*r*h)\3 frist find the heig\[h=\sqrt{15^{2}-9^{2}} =12\]\[ V=\frac{12\times \pi \times9^{2}}{3}=1.018cm\]